OpenStudy (devonhoward15):

Determine if the expression is a perfect square and if so factor it. a^2-8a+16

OpenStudy (pratyush5):

It is :) It is (a-4)^2

OpenStudy (devonhoward15):

thats what i got but its not there... but there is a (a-4)(a-4)

OpenStudy (devonhoward15):

arent they the same?

OpenStudy (pratyush5):

Yes they are the same

OpenStudy (pratyush5):

because for eg 6^2 = 6*6 = 36

OpenStudy (devonhoward15):

can you help me with 2 more problems? ill tag you

OpenStudy (pratyush5):

Yeah sure